OpenStudy (anonymous):

in a guidebook in italy you find 10 cities you wish to visit but only have time to visit 6 how many unique itineraries are possible for you to visit

OpenStudy (anonymous):

604800?

OpenStudy (anonymous):

not sure what "unique itineraries means, but i think it means order counts. so the answer is \[10\times 9\times 8\times 7\times 6\times 5\]